OpenStudy (idealist10):

How to solve this linear equation: t'-t=1+x

OpenStudy (idealist10):

@ganeshie8 @phi

ganeshie8 (ganeshie8):

find the integrating factor and work the standard method ?

OpenStudy (idealist10):

So the integrating factor is e^-x, right? Can you show the work?

ganeshie8 (ganeshie8):

\[\large (ye^{-x})' = (1+x)e^{-x}\] integrating gives you \[\large ye^{-x} = \int (1+x)e^{-x}~dx\]

ganeshie8 (ganeshie8):

use by parts to evaluate the right side integral

OpenStudy (idealist10):

How did you get the (ye^-x)'?

ganeshie8 (ganeshie8):

Ahh replace y by t it has to be t

ganeshie8 (ganeshie8):

\[\large te^{-x} = \int (1+x)e^{-x}~dx\]

OpenStudy (idealist10):

Yes, I got it.

OpenStudy (idealist10):

Also, do you know the prerequisite of Advanced Engineering Mathematics?

ganeshie8 (ganeshie8):

@SithsAndGiggles

OpenStudy (anonymous):

I wouldn't know, I'm not an engineer...

OpenStudy (idealist10):

That's fine.